# Self Supervised Networks for Learning Latent Space Representations of Human Body Scans and Motions

### v1

@@ 2.3 Learning the Latent Space Geometry
In this section, we specifically focus on the situation where our latent space representation can be decomposed into a Pose Parameter Space (i.e., those parameters that govern the motion/pose of the object) and a Shape Parameter Space (i.e., those parameters that govern the identity/body type of the object). This is a typical feature in the construction of human body latent spaces as is the case, for instance, with the SMPL model. As there is no ground truth for the deformation of one human body into another human body, we will equip the Shape Parameter Space simply with the linear Euclidean geometry. The goal of this section is to learn the geometry of the Pose Parameter Space, so that human motions from 4D data are “geodesics” with respect to this geometry. Namely, our goal is to learn a “geometry” such that the following two operations correspond to natural human motions:
@@
To achieve this goal, we train a model comprised of two maps $f:\mathcal{L}^{F}\to\mathbb{R}^{N}$ and $\pi:\mathbb{R}^{N}\to\mathcal{L}^{F}$. Note, that the dimension of the auxiliary space $\mathbb{R}^{N}$ is usually chosen to be significantly larger that the dimension of the target latent space ${\mathcal{L}}^{F}$, i.e., in our experiments we picked $N=4n$.
Given a 4D sequence $\alpha_{1},...,\alpha_{T}$ of latent codes representing a human motion, we then define the loss function
$\operatorname{Loss}^{\operatorname{MoGen}}(\alpha_{1},...,\alpha_{T})=\frac{1}{2T}\sum_{i=1}^{T}\underbrace{\operatorname{MSE}(\hat{\alpha}_{i},\alpha_{i})}_{\text{Interpolation Loss}}+\underbrace{\operatorname{MSE}(\alpha_{i}^{\prime},\alpha_{i})}_{\text{Extrapolation Loss}}$ (4)
@@
$\hat{\alpha}_{i}=\pi\left(f(\alpha_{0})+i\frac{f(\alpha_{T})-f(\alpha_{0})}{T-1}\right)\quad\text{ and }\quad\alpha_{i}^{\prime}=\pi(f(\alpha_{0})+i(f(\alpha_{1})-f(\alpha_{0}))).$
Here the first term is called Interpolation Loss as it measures the deviation of the linear interpolation based on the first and last time point to the given data to the given human body mini sequence. In contrast, the second term is called the Extrapolation Loss as it measures the deviation of the linear extrapolation based on the first two time points to the given data. It is important to note, that our training ensures $\pi\circ f\approx\operatorname{Id}_{\mathcal{L}}$, but that we will not have $f\circ\pi=\operatorname{Id}_{\mathbb{R}^{N}}$ in general. The extrapolation loss ensures in addition that $\pi$ learns to map elements of $\mathbb{R}^{N}$ which are not in $f(\mathcal{L})$ consistently with neighboring points in $f(\mathcal{L})$. This ensures that our framework learns appropriate behavior near the boundary of feasible human motions. We will also demonstrate the advantage of this feature in the experimental section, cf. Figures 6 and 7. For a schematic description of the general latent space geometry procedure we refer to Figure 3.
Figure 3: Illustration of the proposed MoGeN framework: Given two points in latent space (a point and a deformation direction, resp.), the network $f$ maps them into a high dimensional, lifted latent space, where interpolation (extrapolation, resp.) follow the standard linear paths. Subsequently, they are projected back to the original latent space using the network $\pi$, which allows the user to map them into the space of surfaces using the decoder $F$. The networks $f$ and $\pi$ are trained according to the loss function defined in Equation (4). In the above depiction, the boundary points are plotted in red, whereas the initial deformation direction is plotted in green.
@@ 3 Experimental Results
In this section we will perform several experiments to demonstrate the accuracy of the proposed network models and compare them to the state-of-the-art algorithms in the area. We start by describing the computational infrastructure and datasets used in our experiments.
Computational Infrastructure:
All experiments were performed on a standard home PC with a Intel 3.2 GHz CPU and a GeForce GTX 2070 1620 MHz GPU.
@@
We train our models on the Dynamic FAUST (DFAUST) dataset [5]. This database contains high resolution 4D scans captured at 60 Hz of 10 individuals performing 14 in-place motions. These involve scanning errors and occlusions which lead to to topological noise in the data. These 4D sequences are registered using image texture information and a body motion model creating sequences of meshes with fixed point to point correspondences. Furthermore the meshes in these registered sequences share point to point correspondences with outputs of many latent space models such as [19, 26, 28]. We then split this data so that 80 percent of the sequences are used in training and 20 percent are used in testing. Additionally, for the selected testing sequences we also utilize the corresponding raw scans to evaluate the performance of the proposed VariShaPE model on meshes with topological noise and scanning errors and without fixed point to point correspondence. To train and validate the MoGeN model, we use the VariShaPE architecture to extract latent code representations of the training and testing sequences and subsequently extract mini-sequences of pose parameters from these sequences.
3.0.1 Evaluation methods:
In our experiments, we will evaluate the results using different similarity measures: first we calculate the mean vertex distance, which is only available for registered meshes. Secondly, we evaluate the methods using the varifold metric introduced in Section 2.1. In order to avoid any bias, as our method is partially based on this distance measure, we will also use the Chamfer distance [9, 13] as an additional measure of quality. Note, that only the mean vertex distance and the Chamfer distance have an interpretation in physical units (i.e., in millimeters), but that for the varifold kernel distance such an interpretation is not available.
3.1 VariShaPE: Mapping Registered and Unregistered Scans to SMPL
Registered Unregistered
Chamfer ARAPReg 3D Coded Chamfer ARAPReg 3D Coded
Search VAE [17] [13] Ours Search VAE [17] [13] Ours
Mean Vertex Dist. (mm) 6.3 5.1 2.5 2.1 NA NA NA NA
Varifold Error 0.047 0.038 0.024 0.035 0.048 NA 0.035 0.036
Chamfer Error (mm) 3.1 2.9 2.3 2.5 5.7 NA 3.2 2.9
Time of $10^{3}$ Recon. 6339.4s 2.1s 4403.1s 0.8s 9311.6s NA 7012.3s 1.0s
Training None 2w 12hrs 8hrs None NA 12hrs 8hrs
Table 1: Reconstruction Results: We compute the mean vertex distance for the unseen registered testing data and mean Chamfer and varifold errors for all of the meshes in the registered and unregistered testing sets.
Input Data:
Chamfer Search:
3D Coded [13]:
Ours:
Figure 4: Qualitative Comparison of mesh retrieval methods for unregistered data: We display 6 raw scans of human meshes, reconstructions produced by Chamfer search, reconstructions produced by the 3D-Coded model [13], and reconstructions produced via our model. One can clearly see the superior performance of our model and 3D-Coded as compared to Chamfer search. Investigating the results of 3D-Coded in more detail one can see that the output of this model fails to accurately capture fine (textural) details of the human body shapes, such as clothes or the muscles structure on the core. In contrast our model performs significantly better in regards with these fine details. Additionally, we highlight our model’s robustness to noise and scanning errors, as can be seen in the highlighted box at the very right.
Experimental setup:
In our first set of experiments we aim to demonstrate the capability of the VariShaPE model to learn a map from unregistered meshes into the latent space $\mathcal{L}^{F}$ of the SMPL model. Therefore we choose $G$ as the affine decoder described in [15] and $F$ to be the base, gender-neutral SMPL model [19]. This leads to latent space dimensions of $m=170$ and $n=375$. In the appendix we present an ablation study for different choices of latent space models for both $F$ and $G$ and show that these lead to a comparable performance. The network $\Psi_{\theta}$ is then constructed as a sequence of fully connected layers separated by ReLU activation functions with a total of $3\times 10^{6}$ trainable parameters. We train the model with registered meshes from the DFAUST dataset, i.e., we train the model to minimize the mean square error between the vertices of the training meshes and the reconstruction of the learned parameters of the model. An ablation study on the choice of Loss function is shown in appendix B, where one can see that this simple choice of loss function strikes a good balance between computational complexity and quality of the results.
Comparision methods:
For comparison of our latent code retrieval capabilities we use three other state-of-the-art methods: first we simply run an optimization over the latent space using an L-BFGS optimizer minimizing the Chamfer distance between the target mesh and the reconstruction of the latent code. This method works for both registered and unregistered meshes as the Chamfer distance does not rely on point to point correspondences and we will refer to it as Chamfer search. Second, we compare the resulting reconstructions to the variational auto-encoder network of [17]. This approach is trained with registered data to construct a latent space representation of human bodies which is regularized by an as-rigid-as-possible (ARAP) energy. As this network is constructed for data with fixed point to point correspondences we cannot compare its performance on the unregistered testing set, but only use it for pre-registered data. Finally, we compare our results to those obtained with 3D-Coded [12], which models deformations using an auto-decoder with regularization using again an ARAP energy. In contrast to the auto-encoder network of  [17], 3D-Coded utilizes a PointNet architecture as encoder to enable invariance to parameterization thereby allowing us to apply this method also to unregistered data. We train and validate all of the methods presented in this work with the same training and testing split of the DFAUST dataset as our method.
3.1.1 Results:
A quantitative comparison for the latent code retrieval task is presented in Table 1: as one can see VariShaPE significantly outperforms Chamfer search and the VAE based method [17], while being comparable to 3D-Coded [12]. At the same time VariShaPE has an order of magnitudes lower computational cost than Chamfer search and 3D-coded (approximately 5.000 times faster) and a somewhat lower (but comparable) computational cost than the VAE based approach of [17]. In addition to these quantitative results, we present qualitative results in Figure 4; here we restrict the comparision to Chamfer search and 3D-Coded as we focus the presentation to the more challenging task of retrieving the latent code directly from unregistered and noisy scans. One can clearly see the superior performance of our method and 3D-Coded reflected in these qualitative results. In addition one can observe that our method preforms well even w.r.t. fine details such as the muscle structure on the core, whereas 3D-Coded seems to loose these finer details. Note, that these fine texture-like details are not accurately measured with in the quantitative measurements.
3.1.2 Mesh Invariance of the VariShaPE Model:
Finally, we want to carefully explore the sensitivity of the proposed method under remeshing operations, where we compare our results to those of Chamfer search and 3D-Coded, which, similar to VariShaPE, operate on unregistered data and should be thus insensitive to these operations. Towards this objective, we picked 100 shapes from our testing set in DFAUST and created for each of these human body shapes 100 different representatives sampling vertices from the original mesh and applying meshing operations to the resulting point cloud, see Figure 5 for examples of different representations of the same human body shape. We then compared the obtained reconstruction for all 100 representatives, which should be near identical demonstrating the invariance of the different methods under these resampling operations. A qualitative comparison can be seen in Figure 5, whereas a quantitative analysis is presented in Table 2. As one can see, our method significantly outperforms the competing methods, where the performance issues of the other two methods are distinctive: the first method, 3D-Coded, struggles with producing a consistent reconstruction in the presence of mesh resamplings. The second method, Chamfer search, is more consistent but in general rather mediocre, i.e, there is a small variance in the Chamfer reconstructions but each of them has a relatively large error to the ground truth. In contrast our method produces consistent reconstructions with a small error to the ground truth, thereby demonstrating the robustness w.r.t. remeshing operations.
Figure 5: Qualitative results for mesh invariance: on the left, we display four examples of different meshes representing the same human body shape. On the right, we display the reconstructions of 100 such representations using our model (blue), Chamfer search (green), and 3D-coded (yellow). Our model displays invariance to these remeshing operations by producing consistent and accurate reconstructions of the set of meshes representing the same human body shape.
Chamfer Search 3D Coded [13] Ours
Mean Chamfer Error (mm) 5.7 4.0 2.9
Chamfer Error Variance (mm) 0.086 0.246 0.074
Varifold Error Variance 9.0e-4 2.5e-3 5.7e-4
Vertex Location Variance (mm) 1.25 4.79 .038
Table 2: Quantitative results for mesh invariance: We construct a set 100 classes of 100 reparameterizations of random meshes from our data set. We then present the mean of the intraclass standard deviations for the errors and vertex locations of the reconstructions.
3.2 Motion Interpolation and Extrapolation using MoGeN
Experimental Setup:
Next we investigate the capabilities of our second model, MoGeN, to accurately learn the geometry of a latent space representation of human body motions. We report results for the latent space representation of motions for the base, gender-neutral SMPL model. We take $N=1500$ as dimension for the lifted latent space. The maps $f,\pi$ are constructed as fully connected networks using ReLU activation functions and a combined $2\times 10^{7}$ trainable parameters.
Comparison Methods:
To evaluate the performance of MoGeN we evaluate our approach against two competing state-of-the art methods: first, we compare to linear interpolations in the pose parameter space of the SMPL model, which is a common approach in the SMPL community as linear interpolations correspond to joint rotations. This is a natural baseline to compare our model against and further allows us to visualize differences of the corresponding paths in latent space, c.f Figure 6 and Figure 7. Second, we provide a comparison to ARAPReg, the auto-decoder latent space construction of [17].
Results:
In Table 3 one can observe that the proposed algorithm significantly outperforms the competing methods, which is also visualized in the qualitative examples in Figures 6 and 7. We note that the difference in the quantitative results seems significantly lower than the qualitative results suggest: the reason for this is that large parts of the body remain unmoved in these experiments and thus even the significantly worse results of the competing methods have a decent overlap in large regions of the human body, i.e., in those that do not significantly move.
Finally, we note that the computational times for all three methods are on a similar scale (in the order of 2 milliseconds for each of the methods), while the training time of our method is significantly lower than ARAPReg (12 hours as compared to two weeks).
Input Data:
Linear:
ARAPReg [17]:
Ours:
@@ Results:
Figure 6: Qualitative comparison of MoGeN with linear interpolations in SMPL space and with ARAPReg paths. In red, we display a sequence of meshes from a human motion of the dataset with the boundary points highlighted. In green, we display the path which arises from linear interpolation between the boundary points in the latent space $\mathcal{L}^{F}$ while the yellow sequence shows the results obtained with ARAPReg [17]. Finally, in blue, we display the path produced by the MoGeN interpolation model with the boundary points as inputs. On the right, we visualize these paths in a two-dimensional subspace of $\mathcal{L}^{F}$, with the same color coding (the ARAPReg path is not shown there as it involves a different latent space)
Input Data:
Linear:
ARAPReg [17]:
Ours:
Figure 7: Qualitative comparison of MoGeN with linear extrapolation in SMPL space and based on the ARAPReg framework. In red, we display a body motion sequence extracted from data with the initial two points highlighted. In green is shown the path arising from linear extrapolation from the initial points in the latent space, and in yellow the path obtained from ARAPReg [17]. Finally, in blue, we display the path produced by the MoGeN extrapolation model with the boundary points as inputs. On the right plot, we visualize these paths in a two-dimensional subspace of $\mathcal{L}^{F}$ using the same color coding.
@@
Linear ARAPReg [17] MoGeN (Ours) Linear ARAPReg [17] MoGeN (Ours)
Mean Vertex Error (mm) 1.26 1.03 0.46 1.42 1.27 0.59
Chamfer Error (mm) 0.70 0.66 0.53 0.85 0.82 0.64
Varifold Error 0.0073 0.0069 0.0060 0.0092 0.0088 0.0071
@@
Figure 8: Examples of Motion Transfer: On the right we display the transfer of a motion from data (blue) to the identity of a different target shape (red). Using the VariShaPE model we obtain latent codes for the target motion and shape. We then combine the pose parameters of the motion with the identity parameters of the target to produce an transferred motion (purple). On the left we display the transfer of body motion from data (blue) onto the pose and identity of the target shape (red). Using MoGeN, we obtain lifted pose parameters for the sequence and target shape and apply a simple translation in the lifted space. We map the resulting path back to to the latent space and combine the transferred pose parameters with the target identity parameters producing the transferred motion (purple).
4 Applications to Motion Transfer, Generative Modeling and Interpolation between 4D-data.
In this final section, we complement the results of the previous section by presenting a few different possible applications of the proposed methodology.
4.1 Motion transfer
A first example of application is real-time motion transfer, i.e., we consider the problem of transferring a motion from one human body shape to another, cf. Figure 8 for two different examples. Combining VariShaPE and MoGeN immediately leads to a natural and efficient algorithm for that purpose: given a motion sequence (represented by a series of meshes) and a target shape, we map all those meshes in the SMPL latent space ${\mathcal{L}}^{F}$ using the VariShaPE model. This leads to the motion being represented as a path in latent space and the target shape as a point. Subsequently we employ MoGeN to accurately transport the corresponding path of latent codes to the latent code of the target; namely, the path and the point are lifted to the auxiliary latent space of MoGeN, which is equipped with a simple Euclidean geometry, and thus one can apply a simple translation of the path in this space.
@@ 4.1 Motion transfer
4.2 Generative modeling in the space of human body shapes:
Next, we demonstrate how one can combine MoGeN and VariShaPE to obtain a generative model for human bodies based on estimating an empirical distribution from observed training data. In Figure 9, we show 18 random samples from an empirical distribution constructed using 30,000 scans. Note, that this procedure is only feasible due to the light speed encoding capabilities for raw scans via VariShaPE: for our experiment the encoding of the whole population of unregistered training scans took less than one minute. Note that, comparatively, the construction of a similar empirical distribution (i.e., the encoding of the 30,000 scans) using 3D-Coded would require over 2 days. Rather than directly applying kernel density estimator in the SMPL space, we followed the same procedure as in the motion transfer application and first lifted the data into the auxiliary latent space of MoGeN. The linear geometry of that space allows us to use kernel density estimation in a more principled way while still capturing the non-linear nature of the space of human body motion, cf. Figure 9 for a schematic of this procedure.

### v2

@@ 3.3 Learning the Latent Space Geometry
In this section, we assume that our latent space representation can be decomposed into a Pose Parameter Space (i.e. those parameters that govern the motion/pose of the object) and a Shape Parameter Space (i.e. those parameters that govern the identity/body type of the object). This is a typical feature in the construction of human body latent spaces as is the case, for instance, with the SMPL latent space. As there is no ground truth for the deformation of one human body into another human body available we will equip the Shape Parameter Space simply with the linear Euclidean geometry. The goal of this section is to learn the geometry of the Pose Parameter Space, so that human motions from 4D data are “geodesics” with respect to this geometry. Namely, our goal is to learn a “geometry” such that the following two operations correspond to natural human motions:
@@
To achieve this goal, we train a model comprised of two maps $f:\mathcal{L}\to\mathbb{R}^{N}$ and $\pi:\mathbb{R}^{N}\to\mathcal{L}$. Given a 4D sequence $\alpha_{1},...,\alpha_{T}$ of latent codes representing a human motion, we define the loss function
$L(\alpha_{1},...,\alpha_{T})=\frac{1}{2T}\sum_{i=1}^{T}\underbrace{\operatorname{MSE}(\hat{\alpha}_{i},\alpha_{i})}_{\text{Interpolation Loss}}+\underbrace{\operatorname{MSE}(\alpha_{i}^{\prime},\alpha_{i})}_{\text{Extrapolation Loss}}$
@@
$\hat{\alpha}_{i}=\pi\left(f(\alpha_{0})+i\frac{f(\alpha_{T})-f(\alpha_{0})}{T-1}\right)$
and
$\alpha_{i}^{\prime}=\pi(f(\alpha_{0})+i(f(\alpha_{1})-f(\alpha_{0}))).$
Here, the first term is called Interpolation Loss as it measures the deviation of the linear interpolation based on the first and last time point to the given data (i.e., the given human body mini sequence). On the other hand, the second term is called the Extrapolation Loss as it measures the deviation of the linear extrapolation based on the first two time points of the given latent code sequence. It is important to note that our training ensures $\pi\circ f\approx\operatorname{Id}_{\mathcal{L}}$, but that we will not have $f\circ\pi=\operatorname{Id}_{\mathbb{R}^{N}}$ in general. The extrapolation loss term ensures in addition that $\pi$ learns to map elements of $\mathbb{R}^{N}$ which are not in $f(\mathcal{L})$ consistently with neighboring points in $f(\mathcal{L})$. This ensures that our framework learns appropriate behavior near the boundary of feasible human motions. We will also demonstrate the advantage of this feature in the experimental section, cf. Figure 5. For a schematic description of the general latent space geometry procedure, we refer to Fig. 3.
@@ 4 Experimental Results
Experimental Setup:
All experiments were performed on a standard home PC with a Intel 3.2 GHz CPU and a GeForce GTX 2070 1620 MHz GPU, where we used the following setups for our network architectures: for the VariShaPE model we choose $G$ as the affine decoder described in [5, 6] and $F$ to be the base, gender-neutral SMPL model [1]. In Section 4.4 below, we perform an ablation study for different choices of latent space models for both $F$ and $G$. Consequently, $m=170$ and $n=375$. The network $\Psi_{\theta}$ is then constructed as sequence of fully connected layers separated by ReLU activation functions with a total of $3\times 10^{6}$ trainable parameters.
Furthermore, we report results of the proposed MoGeN model for the latent space representation of motions with respect to the base, gender-neutral SMPL model. As mentioned above, the dimension of this latent space model is $375$. For the dimension of the lifted latent space we take $N=1500$. The maps $f,\pi$ are constructed as fully connected networks using ReLU activation functions and a combined $2\times 10^{7}$ trainable parameters.
Registered Unregistered
Cham. VAE Ours Cham. VAE Ours
Mean Vertex Dist. 0.063 0.051 0.021 NA NA NA
Varifold Error 0.047 0.038 0.035 0.048 NA 0.036
Chamfer Error 0.031 0.029 0.025 0.057 NA 0.029
Time of $10^{3}$ Recon. 6339.4s 2.1s 0.8s 9311.6s NA 1.0s
Training None 2w 8hrs None NA 8hrs
Table 1: Reconstruction Results: We compute the mean vertex distance for the unseen registered testing data and mean Chamfer and varifold errors for all of the meshes in the registered and unregistered testing sets. For the unregistered data we display, on the right panel, the cumulative error distributions for each method and evaluation metric.
Chamfer Search:
Input Data:
Ours:
Figure 4: Qualitative Comparison of SMPL latent code retrieval methods for unregistered data: We display 7 raw scans of human meshes (middle), reconstructions produced by minimizing the Chamfer distance (top), and reconstructions produced via our model (bottom). Additionally, we highlight our model’s robustness to noise and scanning errors, as can be seen in the highlighted box at the very right.
@@
We train our networks on the Dynamic FAUST (DFAUST) dataset [35]. This dataset contains high quality 4D scans captured at 60 Hz of 10 individuals performing 14 in-place motions. Due to the high speed of the recording, these scans have topological noise and scanning errors. These 4D sequences are registered using image texture information and a body motion model creating sequences of meshes with fixed point to point correspondences. Furthermore the meshes in these registered sequences share point to point correspondences with outputs of many latent space models such as [1, 2, 7]. We then split this data so that 80 percent of the sequences are used in training and 20 percent are used in testing. Additionally, for the selected testing sequences we also utilize the corresponding raw scans to evaluate the performance of the proposed VariShaPE model on meshes with topological noise and scanning errors and without fixed point to point correspondence. To train and validate the MoGeN model, we use the VariShaPE architecture to extract latent code representations of the training and testing sequences and subsequently extract mini-sequences of pose parameters from these sequences.
@@
Figure 5: Qualitative comparison of MoGeN with linear interpolations in SMPL space. In red, we display a sequence of meshes corresponding to a human motion from data with the boundary points highlighted. In green, we display the path which arises from linear interpolation between the boundary points in the latent space $\mathcal{L}$. Finally, in blue, we display the path produced by the MoGeN interpolation model with the boundary points as inputs. On the right we visualize these paths in two coordinates of the latent space with the same color coding.
Figure 6: Examples of Motion Transfer: On the right we display the transfer of a motion from data (blue) to the identity of a target shape (red). Using the VariShaPE model we obtain latent codes for the target motion and shape. We then combine the pose parameters of the motion with the identity parameters of the target to produce an transferred motion (purple). On the left we display the transfer of human motions from data (blue) onto the pose and identity of the target shape (red). Using MoGeN, we obtain lifted pose parameters for the sequence and target shape and apply a simple translation in the lifted space. We map the resulting path back to to the latent space and combine the transferred pose parameters with the target identity parameters producing the transferred motion (purple).
@@
Linear [9]-AD MoGeN Linear [9]-AD MoGeN
L2 Vertex Error 0.0126 0.0103 0.0046 0.0142 0.0127 0.0059
Varifold 0.0073 0.0069 0.0060 0.0092 0.0088 0.0071
@@
4.1 Mapping Registered and Unregistered Scans to SMPL
In our first set of experiments we aim to demonstrate the capability of our VariShaPE architecture to learn a map from unregistered meshes into the latent space $\mathcal{L}$ of the SMPL model. As described in the setup section we train the model with registered meshes from the DFAUST dataset. As such, we train the model to minimize the mean square error between the vertices of the training meshes and the reconstruction of the learned parameters of the model. For comparison of our latent code retrieval capabilities we use two other state-of-the-art methods: first we run an optimization [16, 10] over the latent space using an L-BFGS optimizer minimizing the Chamfer distance between the target mesh and the reconstruction of the latent code. This method works for both registered and unregistered meshes as the Chamfer distance does not rely on point to point correspondences. Second, we compare the resulting reconstructions to the variational auto-encoder network of [9]. This approach is trained with registered data to construct a latent space representation of human bodies which is regularized by an as-rigid-as-possible (ARAP) energy. As this network is constructed for data with fixed point to point correspondences we cannot compare its performance on the unregistered testing set, but only use it for pre-registered data. A quantitative comparison for the latent code retrieval task is presented in Tab. 1, where we use three different measures to quantify the quality of a reconstruction: (1) the mean vertex distance (only available for registered meshes); (2) the Chamfer distance and (3) the varifold distance. As one can see VariShaPE significantly outperforms the competing methods, while at the same time having an order of magnitudes lower computational cost then the optimization based methods and a somewhat lower (but comparable) computational cost than the VAE based approach. In addition to these quantitative results, we present qualitative results in Fig. 4; here we restrict the comparison to Chamfer search as we focus the presentation to the more challenging task of retrieving the latent code directly from unregistered and noisy scans. See also Figure 10 in Appendix A for a comparison to VAE.